OpenStudy (pokemon23):

If possible,express the binomial as the difference of the squares of monomials; if not possible,tell why.

OpenStudy (pokemon23):

\[r ^{2} +s ^{2}\]

OpenStudy (pokemon23):

prime binomial?

OpenStudy (pokemon23):

what's a prime binomial unable to be factor right ?

OpenStudy (unklerhaukus):

\[(r+is)(r-is)\]\[=r^2-irs+irs-i^2s\]\[=r^2+s^2\]

OpenStudy (pokemon23):

Gracias